# language=R,texcl=true,style=vs

### v2

@@ Abstract
Energy statistics ($\mathcal{\varepsilon}$-statistics) are functions of distances between statistical observations. This class of functions has enabled the development of non-linear statistical concepts, termed distance variance, distance covariance, distance correlation, etc. The computational burden associated with the $\mathcal{\varepsilon}$-statistical quantities is really heavy and when the data reside in the multivariate space, the task becomes even harder. We alleviate this cost by tremendously reducing the memory requirements and essentially making the computations faster. We show the process for the cases of (univariate and multivariate) distance variance, distance covariance, (partial) distance correlation, energy distance and hypothesis testing for the equality of univariate distributions.
@@ 1 Introduction
The concept of energy statistics ($\mathcal{\varepsilon}$-statistics) was introduced through the works of Székely and Rizzo. In 2007 they introduced the distance correlation (Székely et al.,, 2007) as a measure of non-linear correlation and, broadly speaking of dependence between two random variables in arbitrary dimensions. Even though the two main works of Székely and Rizzo (Székely et al.,, 2007) and (Székely and Rizzo,, 2009) have gathered more than 4,000 citations on google.scholar, one could say that they are widely used. However, the computational burden associated with this quantity could perhaps be seen as an obstacle in the wide-spread applicability of this new correlation, which was highly praised by Speed, (2011) who termed it the correlation of the 21st century.
Huo and Székely, (2016) proposed a fast implementation of the distance correlation, but for the univariate case. Their algorithm possesses a computational complexity $O(nlogn)$, which is the same as required for the computation of the Spearman correlation. Their algorithm allowed for the computation of the distance correlation even with millions of observations. However, for the case of 2 or more dimensions the current implementations in the package energy (Rizzo and Szekely,, 2024) has a limit in the number of observations (or sample size), and the same is true for the distance variance, distance covariance and energy distance.
@@ 2.1 Energy distance
2.2 Testing for equality of two distributions
To test for the equality of two univariate and multivariate distributions, the test statistic is the energy distance (1) and the test statistic becomes $T=\frac{nm}{n+m}\mathcal{\varepsilon}\left({\bf x},{\bf y}\right)$. Permutations are used, where shuffling takes place between the two vectors and Eq. (LABEL:uenergy) is computed. This process is repeated $R$ times, and the proportion of times the energy distance test statistics of the permuted data exceeds the energy distance test statistic computed at the observed data serves as the approximate p-value.
@@ 2.4 Fast and light-weight computation of the distance quantities
The R packages energy (Rizzo and Szekely,, 2024) and Rfast (version 2.1.1) (Papadakis et al.,, 2024) compute the distance (co)variance using Equations (2a) and (2b). The problem with that approach is that one needs to compute the distance matrix and store it, and in the case of distance covariance or distance correlation, two matrices must be computed and stored. If the sample size of the observations is on the order of tens of thousands, the computer must be equipped with a really large memory to perform the required computations. Alternatively, the use may use space from their hard drive, an approach that solves the problem of big memory, but does not fully attack the problem of speed. A different approach would be to compute the lower triangular matrix, which has less memory requirements, but the computational cost will only be slightly reduced.
@@ 2.4.1 Computation of distance variance, covariance, correlation and partial correlation
For the univariate distance variance we can exploit the identity used in the Gini coefficient to compute the second term, $\gamma_{2}$, of Eq. (7)
@@
where $x_{(i)}$ denote the ordered values, in ascending order. Regarding the denominator of $\gamma_{1}$, we can do the following
$\sum_{i\neq j}^{n}a_{ij}^{2}=\sum_{i\neq j}^{n}|x_{i}-x_{j}|^{2}=n\sum_{i=1}^{n}x^{2}-\left(\sum_{i=1}^{n}x\right)^{2}.$
Since the sum of all numbers has already been computed since it is required in $\gamma_{1}$ we do not need to compute it twice and also we having to compute the sum of all pairwise squared differences. Finally, for the partial distance correlation, we have implemented it in R by computing the necessary distance correlations.
@@ 2.4.2 Computation of energy distance
Regarding the energy distance (1) a similar trick was applied. The energy consists of three quantities, the sum of all pairwise distances between the matrix $\bf X$ and $\bf Y$, and the sum of all pairwise distances of $\bf X$ and $\bf Y$. We compute the pairwise distances again ”on-the-fly”, that is, compute each pairwise distance and sum it. This way avoids the computation of the whole distance matrix.
@@ 2.4.3 Computations for testing the equality of two distributions
It was mentioned earlier that random permutations occur repeatedly. This means that the constant factor $\frac{nm}{n+m}$ that multiplies the energy distance $(\varepsilon({\bf x},{\bf y}))$ does not influence the p-value. In the case of univariate distributions we can break Eq. (1) into its three terms
@@
Fortunately, in the univariate case, we can again use the identity used in the Gini coefficient to compute the sum of all pairwise distances, without actually computing the distances
@@
For the case of multivariate distributions we can compute the first term ($\gamma_{1}$) and the other two ($\gamma_{2}$ and $\gamma_{3}$) using the commands dista(x, result = "sum") and Dist(x, result = "sum"), respectively, both available in the Rfast library. The resulting function [^3] See Appendix for an R implementation. is not faster than the available implementation in the energy library, but it is memory-saving. Note that the command Dist(x, result = "sum") can also facilitate the memory-efficient computation of the multivariate normality test (Székely and Rizzo,, 2023) as well.
@@ 4 Measuring the computational cost and speed-up factors
We show the time improvements comparing the functions dcor() and edit() from the packages energy, and Rfast (most recent implementation). The experiments took place using a Dell laptop with Intel Core i5-1053G1 CPU at 1GHz, with 256 GB SSD, 8 GB RAM and Windows installed. Using a range of sample sizes (n=(1000, 2000, 5000, 10000, 20000, 50000)) and dimensions (p=(2, 5, 10, 20)) we generate two random matrices, and compute the running time [^5] The time was measured using the built-in command system.time(). each function requires to compute the aforementioned quantities. Table 1 presents the running time (in seconds) required by each implementation [^6] For the case of sample sizes $n=20000$ and higher, the following message appeared ”Error in .dcov(x, y, index) : ’R_Calloc’ could not allocate memory (20000 of 8 bytes)”..
We repeated the examples of the distance correlation but this time compared to the running time required by the function distcor() that is available in the package dcortools (Edelmann and Fiedler,, 2022) with the option to use the memory save algorithm that according to the authors of the package has a computational complexity of $O(n^{2})$ but requires only $O(n)$ memory.
@@ 4.1 Results
Table 1 and 2 presents the running times of energy and Rfast packages required for the computation of the distance correlation, energy distance, and partial distance correlation, respectively [^7] We emphasize that the implementation of the partial distance correlation is in R.. Table 3 presents the estimated scalability of the Rfast::dcor with regards to the sample size, and figure 1 visualizes Tables 1 and 2 and also contains the speed-up factor of Rfast::dcor compared to dcortools::distcor.
For the distance correlation the speed-up factor [^8] This is the ratio of the running time required by the functions in the package energy divided by the running time required by the functions in the package Rfast. ranges from 5.5 (for the case of $n=1000$ and $p=20$) to 119 (for the case of $n=15000$ and $p=10$) with an average of 33. For the energy distance the speed-up factor ranges from 4.6 (for the case of $n=1000$ and $p=10$) and goes up to 88 (for the case of $n=15000$ and $p=2$) with an average of 26. For either quantity, there are two important conclusions. Finally, for the partial distance correlation, the speed-up factor ranges from 7.33 (for the case of $n=1000$ and $p=20$) up to 156 (for the case of $n=15000$ and $p=2$).
At first, the speed-up factor increases with increasing sample size and second, the implementation in the package energy cannot compute those quantities when the sample size exceeds $15000$ due to large memory requirements. This implies that, given the specifications of the computer used, installation of a larger memory that would allow the quantities to be computed with the energy package, would lead to higher speed-up factors with increasing sample sizes.
@@
Finally, Table 4 contains the running time to perform the test of equality of two univariate distributions where the p-value is computed based on 999 permutations. Evidently, memory limitations forbid the application of the test with tens of thousands of observations, but this time, there is an extra limitation, that of running time. In the case of 15000 observations, the energy::eqdist.etest requires a tremendous amount of time, which is more than a single day. In contrast, the pure R implementation requires 11 seconds for the case of 50000 observations. Hence, the speed-up factors need not be computed for this test.
@@ 5 Conclusions
We showed a simple trick to alleviate the memory requirements and the computational cost associated with the energy distance, and the distance variance, covariance and correlation measures. Our next goal is to transfer the equality of univariate distributions testing code in C++. The same is true for the implementation of the partial distance correlation but in this case we must also reduce the number of calculations so as to make this function faster. We would also like to create more efficient functions of computing the permutation based p-value for more hypothesis testing procedures.
@@
The R version of the command dvar() to compute the distance variance for a vector
dvar <- function(x, bc = FALSE) {
  n <- length(x)
  i <- 1:n
  x <- Rfast::Sort(x)
  sxi <- cumsum(x)
  sxn <- sxi[n]
  ai <- (2 * i - n) * x + sxn - 2 * sxi
  #D <- Rfast::Dist(x, square = TRUE, result = "sum")
  D <- n * sum(x^2) - sxn^2
  if ( bc ) {
    a <- 2 * D / ( n * (n - 3) ) - 2 / ( n * (n - 2) * (n - 3) ) * sum(ai^2) +
         sum(ai)^2 / (n * (n - 1) * (n - 2) * (n - 3) )
  } else  a <- 2 * D/n^2 - 2/n^3 * sum(ai^2) + sum(ai)^2/n^4
  sqrt(a)
The R version of the command pdcor() to compute the partial distance correlation
pdcor <- function(x, y, z) {
  a1 <- Rfast::dcor(x, y, bc = TRUE)\$dcor
  a2 <- Rfast::dcor(x, z, bc = TRUE)\$dcor
  a3 <- Rfast::dcor(y, z, bc = TRUE)\$dcor
  up <- a1 - a2 * a3
  down <- sqrt( 1 - a2^2 ) * sqrt( 1 - a3^2 )
  up / down
The R version of the command eqdist.etest() to test the equality of two univariate or multivariate distributions
@@
      id <- Rfast2::Sample.int(n, ni)
      xp <- Rfast::Sort(z[id])
@@
    ## multivariate
  } else {
    nx <- dim(x)[1]  ;  ny <- dim(y)[1]
    n <- nx + ny
    stat <- Rfast::edist(x, y)
    z <- rbind(x, y)
    for ( k in 1:R ) {
      id <- Rfast2::Sample.int(n, nx)
      pstat[k] <- Rfast::edist(z[id, ], z[-id, ])

### v4

@@ Abstract
Energy statistics, also known as $\mathcal{\varepsilon}$-statistics, are functions of distances between statistical observations. This class of functions has enabled the development of non-linear statistical concepts, such as distance variance, distance covariance, and distance correlation. However, the computational burden associated with $\mathcal{\varepsilon}$-statistics is substantial, particularly when the data reside in multivariate space. To address this challenge, we have developed a method to significantly reduce memory requirements and accelerate computations, thereby facilitating the analysis of large data sets. The following cases are demonstrated: univariate and multivariate distance variance, distance covariance, partial distance correlation, energy distance, and hypothesis testing for the equality of univariate distributions.
@@ 1 Introduction
Székely et al., (2007) pioneered the concept of energy statistics, which they termed ”energy statistics” or $\mathcal{\varepsilon}$-statistics. In their seminal work, they introduced the distance correlation (Székely and Rizzo,, 2009) as a measure of nonlinear correlation and, in essence, dependence between two random variables in arbitrary dimensions. Despite the fact that the two primary works of Székely et al., (2007), Székely and Rizzo, (2009) have amassed over 4,000 citations on Google Scholar, it could be argued that they are widely employed. However, the computational burden associated with this quantity could be perceived as a hindrance to the widespread applicability of this novel correlation, which was highly lauded by Speed, (2011) who designated it as the correlation of the 21st century.
Distance correlation has been successfully used in many fields, such as variable selection (Li et al.,, 2012), network analysis (MacCarron et al.,, 2023), and time series (Edelmann et al.,, 2019). Its uses extend beyond statistics, such as in finance (Ugwu et al.,, 2023), bioinformatics (Hou et al.,, 2022), and physics (Kasieczka and Shih,, 2020) to name a few.
Huo and Székely, (2016) proposed a fast implementation of the distance correlation, but for the univariate case. Their algorithm possesses a computational complexity class $O(nlogn)$, which is the same as required for the computation of the Spearman correlation. Their algorithm allowed for the computation of the distance correlation even with millions of observations.However, for the case of two or more dimensions, the current implementations in the package R package energy (Rizzo and Szekely,, 2024). The latter is limited by the number of observations (or sample size), and this limitation also applies to the distance variance, distance covariance, and energy distance. However, the partial distance correlation, the energy distance, and the equality of distributions tests that rely on the latter are still computationally expensive and require substantial memory when the number of observations is measured in tens of thousands or more.
The present paper makes two contributions. Firstly, it demonstrates methods for reducing the computational burden and memory requirements of certain energy statistics quantities. Secondly, it elucidates techniques for increasing the speed of calculations while reducing memory requirements. In the multivariate case, the paper employs the formulas presented in Székely and Rizzo, (2023). In the univariate case, it utilizes mathematical identities. These approaches are instrumental in expediting the calculation of energy statistics, including the energy distance, a metric of distribution equality that hinges on energy distance, distance variance, covariance, correlation, and partial distance correlation.
The next section briefly describes some quantities from the energy statistics and shows how to make computationally efficient, while Section 3 presents the relevant functions in the R package Rfast (Papadakis et al.,, 2025). Section 4 compares the Rfast implementations to implementations in other R packages and finally Section 5 concludes the paper.
@@ 2.2 Distance variance, covariance and correlation
2.3 Testing for equality of two distributions
To test for the equality of two univariate (and multivariate distributions), the test statistic is the energy distance (1) and the test statistic becomes $T=\frac{nm}{n+m}\mathcal{\varepsilon}\left({\bf x},{\bf y}\right)$. Permutations are used, where shuffling takes place between the two vectors and Eq. (1) is computed. This process is repeated $R$ times, and the proportion of times the energy distance test statistics of the permuted data exceeds the energy distance test statistic computed at the observed data serves as the approximate p-value.
@@ 2.4 Fast and light-weight computation of the distance quantities
The R packages energy (Rizzo and Szekely,, 2024) and Rfast compute the distance (co)variance using Equations (2a) and (2b), respectively. However, this approach is problematic because it necessitates the computation and storage of the distance matrix. In the case of distance covariance or distance correlation, two matrices must be computed and stored. In scenarios where the sample size of observations approaches tens of thousands, it is imperative that the computer system possesses substantial memory capacity to facilitate the requisite computations. Alternatively, one may utilize available storage space on the hard drive, though this approach only addresses the memory constraints but does not fully address the issue of computational speed.A different approach involves computing the lower triangular matrix, which requires less memory, but the resulting reduction in computational cost is only marginal.
@@ 2.4.1 Computation of distance variance, covariance, correlation and partial correlation
For the univariate distance variance, the identity in the Gini coefficient facilitates fast computation of the second term $\gamma_{2}$, of Eq. (7),
@@
where $x_{(i)}$ denote the ordered values, in ascending order. Regarding the numerator of $\gamma_{1}$, we can do the following
$\sum_{i\neq j}^{n}a_{ij}^{2}=\sum_{i\neq j}^{n}|x_{i}-x_{j}|^{2}=n\sum_{i=1}^{n}x_{i}^{2}-\left(\sum_{i=1}^{n}x_{i}\right)^{2}.$
Since the sum of all numbers has already been computed (it is required in $\gamma_{1}$) there is no need to compute it twice, thus avoiding the computation of the sum of all pairwise squared differences. Finally, for the partial distance correlation, the implementation in R computes the necessary distance correlations and is thus computationally efficient.
@@ 2.4.2 Computation of energy distance
The energy distance (1) exploits a similar trick. The energy consists of three quantities, the sum of all pairwise distances between the matrix $\bf X$ and $\bf Y$, and the sum of all pairwise distances of $\bf X$ and $\bf Y$. The pairwise distances are computed again ”on-the-fly”, that is, each pairwise distance is computed and summed, avoiding the computation of the whole distance matrix.
@@ 2.4.3 Computations for testing the equality of two distributions
As mentioned earlier, random permutations occur repeatedly. Thus, the constant factor $\frac{nm}{n+m}$ that multiplies the energy distance $(\varepsilon({\bf x},{\bf y}))$ does not influence the p-value. With univariate distributions Eq. (1) consists of three terms
@@
The use of the identity in the Gini coefficient ensures fast computation the sum of all pairwise distances,
@@
For the case of multivariate distributions, the commands dista(x, result = "sum") and Dist(x, result = "sum") (both available in the Rfast library) compute the first term ($\gamma_{1}$) and the other two ($\gamma_{2}$ and $\gamma_{3}$), respectively. The resulting function [^3] See Appendix for an R implementation. is not faster than the available implementation in the energy library, but it is memory-saving. Note that the command Dist(x, result = "sum") can also facilitate the memory-efficient computation of the multivariate normality test (Székely and Rizzo,, 2023) as well.
Two additional tricks are a) the fact that the combined samples are sorted and b) that the total sum does not change. When permuting the data, one needs to sort the observations that fall into the first permuted sample, the rest of the observations that form the second permuted sample are already sorted. Also, instead of sorting the observations of the first permuted sample, sort their indices, since it is faster to sort integers than numeric numbers. The $\gamma_{2}$ quantity involves calculating the sum of each permuted sample. Since the total sum does not change, one only needs to compute the sum of the observations that fall within the permuted sample, the sum of the observations of the second permuted sample is easily obtained by subtraction.
@@ 4 Measuring the computational cost and speed-up factors
This section illustrates the time improvements comparing the functions dcor() and edit() from the packages energy, and Rfast (most recent implementation). The experiments were conducted utilising a Dell laptop equipped with Intel Core i5-1053G1 CPU at 1GHz, with 256 GB SSD, 8 GB RAM and Windows installed. Using a range of sample sizes (n=(1000, 2000, 5000, 10000, 20000, 50000)) and dimensions (p=(2, 5, 10, 20)) we generate two random matrices, and compute the running time [^5] The time was measured using the built-in command system.time(). each function requires to compute the aforementioned quantities. Table 1 presents the running time (in seconds) required by each implementation [^6] For the case of sample sizes $n=20000$ and higher, the following message appeared ”Error in .dcov(x, y, index) : ’R_Calloc’ could not allocate memory (20000 of 8 bytes)”..
The distance correlation examples were repeated, but this time compared the running time required by the function distcor() that is available in the package dcortools (Edelmann and Fiedler,, 2022) with the option to use the memory save algorithm that according to the authors of the package has a computational complexity of $O(n^{2})$ but requires only $O(n)$ memory.
@@ 4.1 Results
Table 1 and 2 presents the running times the energy and Rfast packages require for the computation of the distance correlation, energy distance, and partial distance correlation, respectively [^7] We emphasize that the implementation of the partial distance correlation is in R.. Table 3 presents the estimated scalability of the Rfast::dcor with regards to the sample size, and figure 1 visualizes Tables 1 and 2 and also contains the speed-up factor of Rfast::dcor compared to dcortools::distcor.
For the distance correlation, the speed-up factor [^8] This is the ratio of the running time required by the functions in the package energy divided by the running time required by the functions in the package Rfast. ranges from 5.5 (for the case of $n=1000$ and $p=20$) to 119 (for the case of $n=15000$ and $p=10$) with an average of 33. For the energy distance, the speed-up factor ranges from 4.6 (for the case of $n=1000$ and $p=10$) and goes up to 88 (for the case of $n=15000$ and $p=2$) with an average of 26. Finally, for the partial distance correlation, the speed-up factor ranges from 7.33 (for the case of $n=1000$ and $p=20$) up to 156 (for the case of $n=15000$ and $p=2$).
For either quantity, there are two important conclusions. The speed-up factor increases with increasing sample size and second, the implementation in the package energy cannot compute those quantities when the sample size exceeds $15000$ due to large memory requirements. This implies that, given the specifications of the computer used, installation of a larger memory that would allow the quantities to be computed with the energy package, would lead to higher speed-up factors with increasing sample sizes.
@@
Finally, Table 4 contains the running time to perform the test of equality of two univariate distributions when the p-value is computed based on 999 permutations. Evidently, memory limitations forbid the application of the test with tens of thousands of observations, but this time, there is an extra limitation, that of running time. In the case of 15,000 observations, the energy::eqdist.etest requires a tremendous amount of time, which is more than a single day. In contrast, the pure R implementation requires 11 seconds for the case of 50000 observations. Hence, the speed-up factors need not be computed for this test.
@@ 5 Conclusions
We showed a simple trick to reduce the memory and computational cost associated with the energy distance and the distance variance, covariance, and correlation measures. Our next goal is to transfer the code for testing the equality of univariate distributions to C++. The same is true for the partial distance correlation implementation, but in this case we also need to reduce the number of computations to make this function faster. We would also like to create more efficient functions for computing the permutation-based p-value for more hypothesis testing procedures.
@@
The R version of the command eqdist.etest() to test the equality of two univariate
@@
      id <- Rfast::Sort.int( Rfast2::Sample.int(n, ni) )
      xp <- z[id]